OpenStudy (anonymous):

Please help me with an addition formula problem! Solve the system by the elimination method. Check your work. 2a + 3b = 6 5a + 2b - 4 = 0

OpenStudy (anonymous):

Okay so your first step is to solve for one of the variable in one of the equations.

OpenStudy (anonymous):

Then take whatever is on the other side of the equation and substitute it for the variable in the other equation

OpenStudy (anonymous):

But I'm not sure how to do that.

OpenStudy (anonymous):

Is it 2a + 3b = 6 6ab = 6 divide both sides by 6 ab = 1 ???

OpenStudy (baldymcgee6):

@wio is reffering to the substitution method, the question requires the use of elimination.

OpenStudy (anonymous):

Okay, I'm confused now.

OpenStudy (baldymcgee6):

Substitution and elimination are two different methods to solve systems of equations, which one do you want to use?

OpenStudy (anonymous):

Let's use elimination.

OpenStudy (baldymcgee6):

2a + 3b = 6 5a + 2b - 4 = 0 By examining the two equations, we try and look to do something to one or both of them to make two of the values the same, so for example, if we multiply equation one by a factor of 5 and multiply eqn two by a factor of 2, we will get: 10a + 15b = 30 10a + 4b - 8 = 0 Notice how we have 10 in both of them

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

So can we just take out the 10 and leave a?

OpenStudy (baldymcgee6):

no... if we subtract the two equations, then we will get ride of a completely. 10a + 15b -30 = 0 - 10a + 4b - 8 = 0 -------------------- 11b - 38 = 0

OpenStudy (anonymous):

oh, ok

OpenStudy (baldymcgee6):

then we can solve for b right there. 11b - 38 = 0 11b = 38 b= 38/11 Sub that b value back into either of the equations to find a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2a + 3b = 6....(1) 5a + 2b - 4 = 0 ...(2) (2a + 3b = 6) * 2...(3) (5a + 2b - 4 = 0) * 3...(4) elimation: (3) - (4) a = ? then substitude a = ? into (3) or (4) to find b.

OpenStudy (anonymous):

ok, so a - 4

OpenStudy (anonymous):

but 0 times any number is 0, so a = 0?

OpenStudy (anonymous):

@baldymcgee6 is the answer 38/11,0?

OpenStudy (baldymcgee6):

do those answers solve the system?

OpenStudy (anonymous):

No, they don't

OpenStudy (anonymous):

but that's all I could come up with.

OpenStudy (anonymous):

What do I do?

OpenStudy (baldymcgee6):

We must have done something wrong.. let me check

OpenStudy (baldymcgee6):

2a + 3b = 6 5a + 2b - 4 = 0 10a + 15b -30 = 0 - 10a +4b - 8 = 0 ---------------------- 11b -22 = 0 11b = 22 b = 22/11 = 2

OpenStudy (baldymcgee6):

my bad, sorry about that. Just a simply subtraction error there.

OpenStudy (anonymous):

I think I've tried 2,0, but that didn't work..

OpenStudy (anonymous):

@baldymcgee6 2,0 didn't work, should I try 0,2?

OpenStudy (baldymcgee6):

2a + 3b = 6 5a + 2b - 4 = 0 if b =2 and a = 0 2(0) + 3(2) = 6 TRUE 5(0) + 2(2) - 4 = 0 TRUE

OpenStudy (anonymous):

Okay, so 6,0?

OpenStudy (baldymcgee6):

no, our answer is correct, b =2 a = 0

OpenStudy (anonymous):

do I put it in the order of 2,0, or 0,2?

OpenStudy (baldymcgee6):

they are not co-ordinate points, a=0, b=2

OpenStudy (anonymous):

Well, that's indeed strange that it counted as wrong. I'll ask my teacher!

OpenStudy (baldymcgee6):

okay

OpenStudy (anonymous):

@baldymcgee6 hey you got it right guys it was 0,2